OpenStudy (anonymous):

which equation represents the line passing through (6,-3) and (-4,-9)

OpenStudy (anonymous):

which graph? Cartesian?

OpenStudy (anonymous):

it just gives me points

OpenStudy (kainui):

if you are deciding between which graphs, you can plug in the coordinates and see if you get something true out. for instance if you want to see if the point (2,4) falls on the line y=2x then you plug in 2 for x and 4 for y and you get 4=2*2 which is true, so then it must fall on the line. try plugging in (3,5) and you'll see that 5=2*3 is not true, so the point doesn't fall on the line.

OpenStudy (anonymous):

so would it be y+-3x-2, y+3x+2, y+4= -3(x-2), or y+8= -3(x-2)? @Kainui

OpenStudy (kainui):

Plug them in like I said and find out for yourself. Good luck.

OpenStudy (anonymous):

but there aren't any graphs @Kainui

OpenStudy (kainui):

There doesn't need to be, but you can make one from the information you're given. You know what an ordered pair is?

OpenStudy (anonymous):

yes @Kainui

OpenStudy (kainui):

what's an ordered pair, explain it. what's the first and second number mean?

OpenStudy (kainui):

y+4= -3(x-2) replace the x and y in that with the x and y in one of your ordered pair to check if the point falls on this line. I just picked this one randomly, and I have absolutely no idea if this is the right one or not, but I'll help you check it if you try. If not, then you're on your own.